OpenStudy (anonymous):

Integrate (x + sinx)^2 dx

OpenStudy (raden):

u need expand it before integrate

OpenStudy (raden):

(x + sinx)^2 = x^2 + 2xsinx + sin^2 x

OpenStudy (anonymous):

i got upto that part but what am I supposed to do next ?

OpenStudy (raden):

just integrate each term

OpenStudy (raden):

int x^2 dx = 1/3 x^3 , right :)

OpenStudy (anonymous):

first term (x^2) is easy... second term is integration by parts... third term use: sin^2x = 1/2(1-cos2x)

OpenStudy (raden):

int 2xsinx dx ---> use int by part OR use the table integral

OpenStudy (anonymous):

how would u do it by parts? i'm confused with that ... I get it the easy way but not with this method

OpenStudy (raden):

int (2x sinx) dx =